A basket contains 5 white balls and 2 green balls. If three balls are taken from the basket one after another, determine the possible values of the random variable G representing the number of green balls. List the sample space. Determine the probability of the random variable G. *


1
Expert's answer
2022-03-05T07:11:45-0500

Sample space: {WWW, WWG, WGW, GWW, WGG, GWG, GGW}.

"P(G=0)=\frac{1}{7}."

"P(G=1)=\frac{3}{7}."

"P(G=2)=\frac{3}{7}."